# On the Wasserstein distance

### v2

@@ Abstract
In this paper, we revisit the Dobrushin uniqueness theorem for Gibbs measures of lattice systems of interacting particles at thermal equilibrium. In a nutshell, Dobrushin’s uniqueness theorem provides a practical way to derive sufficient conditions on the inverse temperature and/or model parameters assuring uniqueness of Gibbs measures by reducing the uniqueness problem to a suitable estimate of the Wasserstein distance between pairs of 1-point Gibbs measures with different boundary conditions. After proving a general result of completeness for the Wasserstein distance, we reformulate the Dobrushin uniqueness theorem in a convenient form for lattice systems of interacting particles described by Hamiltonians that are not necessarily translation-invariant with possibly infinite-range pair-potentials and with general complete metric spaces as single-spin spaces. Our reformulation includes existence, and covers both classical and the Euclidean version of quantum lattice systems. Further, a generalization to the Dobrushin-Shlosman theorem for translation-invariant lattice systems is given. Subsequently, we give a series of applications of these uniqueness criteria to some high-temperature classical lattice systems including the Heisenberg, Potts and Ising models. An application to classical lattice systems for which the local Gibbs measures are convex perturbations of Gaussian measures is also given.
@@
The existence problem. This may be solved by constructing Gibbs measures as thermodynamic limits. This may be achieved by proving that the family of local Gibbs measures, indexed by an increasing sequence of bounded regions filling the whole lattice, has at least one limit point in the weak topology, and that this limit point is a Gibbs measure. Sufficient conditions were derived by Dobrushin [19, 20, 22] and [23, Thm. 1]. See also [75, Thm. 1.3]. Dobrushin’s existence criteria have been applied to some classical lattice models of Euclidean lattice field theories with single-spin space $\mathbb{R}$ in [17, 11]. Therein, some restrictions on the boundary conditions are necessary. One of the key methods is Ruelle’s technique of superstability estimates in [68, 69, 50] which requires the interactions to be translation invariant, superstable and with pair-potentials growing at most quadratically. Applied to some models of Euclidean lattice field theories with single-spin space $\mathbb{R}$, it is proved in [17] that the family of local Gibbs measures associated with a wide class of boundary conditions has at least one accumulation point in the set of ’superstable’ Gibbs measures, a subset of the set of tempered Gibbs measures, see [17, Thm. 1.2]. This result was extended to some quantum anharmonic lattice systems with superstable interactions in [60, Thm 2.6] by extending Ruelle’s technique to quantum statistical mechanics, see [57, 58]. See also [46, Thm. 3.1] covering a wide class of quantum anharmonic lattice systems. In general, for quantum lattice systems where the single-spin spaces are infinite-dimensional, verifying Dobrushin’s existence criteria turns out to be challenging, see, e.g., [75, Sec. I.5]. A brief review of methods can be found in [7, Sec. 2].
@@
We mention that Dobrushin’s criterion has been applied to some classical lattice models of Euclidean lattice field theories with single-spin space $\mathbb{R}$ in [17, Sec. 2]. The key ingredient is an expression of the Wasserstein distance for probability measures on $\mathbb{R}$ in terms of their distribution functions. The same criterion has also been applied to some quantum anharmonic lattice models of quantum crystals in [3, 5]. See also [46, Thm 3.4]. Therein, the main ingredient is the representation of the Wasserstein distance by the Kantorovich-Rubinstein duality theorem, see, e.g., [30]. For completeness, we mention some other techniques to prove uniqueness of Gibbs measures. Cluster expansion techniques have been used in [61] to prove uniqueness in the high-temperature regime for a class of quantum anharmonic lattice systems with superstable interactions. Note that in 1-dimensional lattices, uniqueness holds at all temperatures. See also [2, 64]. In [4], a quantum crystal with double-well potential is shown to have a phase transition using a Peierls-type argument.
@@
In high-temperature translation-invariant lattice systems with interactions in the Dobrushin uniqueness region, Dobrushin’s uniqueness theorem assures the non-existence of first-order phase transitions. Gross showed in fact the absence of second-order phase transitions. More precisely, Gross proved in [36] that the pressure is twice continuously Gateaux differentiable in the Dobrushin uniqueness region (differentiability is in parameters occurring linearly in the interactions). Prakash extended the Gateaux differentiability to any order in a subset of the Dobruhin uniqueness region, see [63]. Turning to complex analyticity, Gallavotti and Miracle-Sole showed in the case of single-site spin space consisting of two points, that the pressure is analytic at high temperature for a broader class of potentials than that to which the Dobrushin uniqueness theorem applies, see [32]. The techniques used seem to be limited to spin spaces having a finite number of points. For some results on high-temperature analyticity with continuous spin spaces and a general class of many-body long-range interactions, see, e.g., [43]. Herein, Dobrushin’ uniqueness techniques are used. See also [26, 27]. For some general results on analyticity via high-temperature expansions, see, e.g., [29, 42] and references therein.
@@
In this paper, we revisit Dobrushin’s uniqueness theorem of Gibbs measures for lattice systems at thermal equilibrium that are not necessarily translation-invariant, consisting of interacting particles (with possibly long-range interactions) and with a general complete metric space as single-spin space. Both the existence and uniqueness problems are proved in this general setting, along with the Dobrushin-Shlosman generalization of Dobrushin’s uniqueness theorem for translation-invariant lattice systems. Our formulation of the uniqueness theorems cover both classical lattice systems and the Euclidean version of quantum lattice systems.
@@
Our first series of applications include the classical Heisenberg model, $q$-state Potts model and Ising model. With a view to applying the Dobrushin and Dobrushin-Shlosman criteria, our focus is on different techniques to estimate the Wasserstein distance between two probability measures. We start by proving a general result of uniqueness at high temperatures with direct application to the classical Heisenberg model, see Proposition 6.1. Our estimation of the Wasserstein distance is based on an approximation procedure for probability measures. Subsequently, we derive an explicit lower bound for the critical inverse temperature for the $q$-state Potts models in dimension $d=2,3$ (asymptotic for large $q$), followed by the $d$-dimensional Ising model. The results are based on an explicit formula of the Wasserstein distance for probability measures on $q$-point spaces, see Lemma 6.3. Finally, the effectiveness of the Dobrushin-Shlosman criteria is illustrated with the square lattice Ising model in dimension $d=2,3$. The estimation of the Wasserstein distance is based on linear programming techniques, see, e.g., [33] and Section 6.3.1 for a brief overview. This algorithm is used to numerically compute a lower bound for the critical inverse temperature.
@@
Our second series of applications focus on convex perturbations of the Gaussian free field model. After applying the Dobrushin criterion to the Gaussian free field model with $d$-dimensional spin, the uniqueness result is extended when considering a perturbation by a convex self-interaction. Proposition 7.5 covers the case of 1-dimensonial spin while Proposition 7.9 covers the case of $d$-dimensional spins for a particular class of convex potentials. We show that uniqueness holds at all temperatures by constructing suitable couplings with the 1-point Gibbs measures as marginals.
@@
Our paper is organized as follows. In Section 2, we recall some properties of the Wasserstein distance and give the completeness result in Theorem 2.1. After recalling some definitions of local and limit Gibbs distributions, Theorem 3.5 in Section 3 is our reformulation of Dobrushin’s uniqueness theorem for general lattice systems of interacting particles when the single-spin space is a general metric space. This covers classical and Euclidean version of quantum lattice systems. Two corollaries related to the decay of correlations in the case of nearest-neighbour interactions are also given. The proof of Theorem 3.5 and proofs of the corollaries are given in Section 4. Theorem 5.1 in Section 5 is our reformulation of Dobrushin-Shlosman’s uniqueness theorem for translation-invariant lattice systems. In Section 6, we focus on a first series of applications of the uniqueness criteria to classical lattice systems at high-temperatures. Dobrushin’s criterion is applied to the classical Heisenberg model in Section 6.1, the $q$-state Potts model in Section 6.2.1 and the $d$-dimensional Ising model in Section 6.2.3. Before applying the Dobrushin-Shlosman criterion to the 2 and 3-dimensional square lattice Ising model in Section 6.3.2, we recall some results from linear programming techniques in Section 6.3.1 including the algorithm used for the numerical estimations. In Section 7, we focus on a second series of applications to classical lattice systems with convex perturbations. Dobrushin’s criterion is applied to the Gaussian free field model with $n$-dimensional spins in Section 7.1. The perturbation by a convex self-interaction potential is then treated in Section 7.2, see Propositions 7.5 and 7.9. Section 8 contains the proof of Theorem 2.1. The key result for the proof of Theorem 2.1 $(ii)$ is Proposition 8.5. Section 9 is our Appendix. The disintegration theorem for measures is recalled in Section 9.1. The dual problem corresponding to the linear programming problem is discussed in Section 9.2. By way of application, a proof of the Kantorovich-Rubinstein duality theorem is given in Section 9.2.3. The quantum harmonic crystal lattice model is revisited in Section 9.3, see Proposition 9.9.
@@
We now turn to the definition of limit Gibbs distributions. We refer the reader to the disintegration theorem in Sec. 9 for the existence of conditional probability measures.
@@
The conditional probability measure in $(ii)$ above is relative to the projection map $\Pi_{\Lambda^{c}}$, see Sec. 9. Denoting it by $\mu^{\beta}(\cdot\,|\underline{\eta}_{\Lambda^{c}})$ for a fixed configuration $\underline{\eta}_{\Lambda^{c}}\in\mathcal{X}^{\Lambda^{c}}$, $(ii)$ reads
@@
In Theorem 3.5, $(\mathcal{X},\rho)$ is a general complete metric space. The latter could be the infinite-dimensional space of all continuous and periodic functions on $[0,\beta]$ equipped with the supremum norm, so that Theorem 3.5 also applies to quantum lattice systems. See Sect. 9.3.
@@
Note that here we used (4.12). Clearly, the first term in the r.h.s. of (4.17) tends to 0 in the limit $p\rightarrow\infty$. To conclude the proof of (4.14), we use the monotone convergence theorem to take $n\rightarrow\infty$.
@@ 5 A generalization: The Dobrushin-Shlosman Theorem
The constant $r$ in (5.1)-(5.3) may depend on the boundary condition $\eta_{\Lambda_{0}^{c}}$.
@@
Assume now that $\epsilon<\frac{\alpha}{2}$, and choose $\Lambda^{\prime}\subset\mathbb{Z}^{d}$ such that $\Lambda+\Lambda_{0}\subset\Lambda^{\prime}$ and so large that
$R\sum_{k\in\Lambda^{\prime c}}\mathrm{e}^{-\frac{\alpha}{2}|k|}<\frac{1}{2}(1-\lambda).$
@@
From (5.10) along with (5.11) and (5.12), we eventually have
@@
The proof easily extends to tempered Gibbs measures. Alternatively, the exponential decay in (5.1) can be replaced by a sufficiently fast polynomial decay if the distributions satisfy the condition in (3.11).
@@
Assume that $\mathcal{X}$ is either a compact convex subset of a normed space with metric $\rho$, or a compact Riemannian manifold. In the latter case, assume that the metric on $\mathcal{X}$ is given by $\rho(x,y)=\int_{0}^{1}|\zeta(s)|\mathrm{d}s$, where $\zeta$ is a shortest geodesic from x to y. Consider the following formal Hamiltonian with nearest-neighbour interactions defined on $\mathcal{X}^{\mathbb{Z}^{d}}$ as
@@
Further, let $\mu_{0}\in\mathcal{P}(\mathcal{X})$ be a probability measure and, for a given $\beta>0$, let $\mu_{j}^{\beta}=\mu^{\beta}$ denote the associated 1-point Gibbs distribution generated by (6.1). Let $C_{1}>0$ be defined as
@@
In fact, considering $\mathbb{S}^{r}$ as a subset of $\mathbb{R}^{r+1}$, the estimate can be slightly improved to $\beta<(24dJ)^{-1}$. Using the symmetry as in the Ising model in Sec. 6.2 below, the estimate can be improved further.
@@
followed by assumption $(ii)$, we obtain, under conditions (6.7), that
@@
$\mu^{\beta}(A|\underline{\eta}^{\prime})-\mu^{\beta}(A|\underline{\eta})=\\ \frac{1}{Z^{\beta}(\underline{\eta})}\int_{A}\left(\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta}^{\prime})-\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta})\right)\mu_{0}(d\xi_{0})+\frac{\mu^{\beta}(A|\underline{\eta}^{\prime})}{Z^{\beta}(\underline{\eta})}\int_{\mathcal{X}}\left(\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta}^{\prime})-\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta})\right)\mu_{0}(d\xi_{0}).$
@@
Note that $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X}\times F)=\mu^{\beta}(F|\underline{\eta}^{\prime})$ and $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(E\times\mathcal{X})=\mu^{\beta}(E|\underline{\eta})$ so that $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}\in\Xi_{\mathcal{X}}(\mu^{\beta}(\cdot\,|\underline{\eta}),\mu^{\beta}(\cdot\,|\underline{\eta}^{\prime}))$. (This follows from the fact that $\nu_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X})=\tilde{\nu}_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X})=\frac{2\kappa_{\beta}}{1+\kappa_{\beta}}$.) Now,
@@
Theorem 6.1 does not apply to the $q$-state Potts model or the Ising model.
@@
The $q$-state Potts model is a generalization of the Ising model in that the spin values can be any arbitrary integer in $\{1,\ldots,q\}$, see, e.g., [62]. We refer to, e.g., [10, 39], for some exact results in dimension $d=2$.
@@
Below, this expression is evaluated numerically in dimensions $d=2$ and $d=3$. When $d=2$, there are only 8 distinct cases to be considered. Since the interaction is invariant under renaming of spins, the distinct configurations $\underline{s}$ are
$(1,1,1,1),\,\,\,(1,1,1,2),\,\,\,(1,1,2,2),\,\,\,(1,1,2,3),\,\,\,(1,2,3,4).$
Changing one spin in the first, it can only combine with the second as $\underline{s}^{\prime}$. The second can also combine with the third and fourth and also with $(1,1,1,3)$. The third can also combine with the fourth, and the fourth can also combine with the fifth as well as with $(1,3,2,3)$. We therefore have the following pairings:
@@
$\underline{s}$ $=(1,1,2,3),\quad\underline{s}^{\prime}=(1,2,3,3)$
$\underline{s}$ $=(1,1,2,3),\quad\underline{s}^{\prime}=(1,2,3,4).$
The figures below show the graphs of $\rho_{W}$ as a function of $\beta$ for the most relevant of these cases and $q=4$ respectively $q=30$.
@@
In Figures 1 and 2, the line $y=\frac{1}{4}$ has also been drawn, which is the upper bound on the right-hand side of the Dobrushin inequality (3.15). For $q=4$, the curve which intersects the horizontal line in the left-most point corresponds to the case $\underline{s}=(1,1,1,2)$ and $\underline{s}^{\prime}=(1,1,2,2)$. For $q=30$, this curve is overtaken by the case $\underline{s}=(1,1,1,1)$ and $\underline{s}^{\prime}=(1,1,1,2)$. Note that the corresponding curve for $q=4$ is the blue curve with a clear maximum. The change-over occurs at $q=20$, as is clearly visible in the following graph of the Dobrushin lower bound for $\beta_{c}J$ ($\beta_{c}$ is the critical inverse temperature) as a function of $q$:

### v4

@@ Abstract
In this paper, we revisit the Dobrushin uniqueness theorem for Gibbs measures of lattice systems of interacting particles at thermal equilibrium. In a nutshell, Dobrushin’s uniqueness theorem provides a practical way to derive sufficient conditions on the inverse temperature and/or model parameters assuring uniqueness of Gibbs measures by reducing the uniqueness problem to a suitable estimate of the Wasserstein distance between pairs of 1-point Gibbs measures with different boundary conditions. After proving a general result of completeness for the Wasserstein distance, we reformulate the Dobrushin uniqueness theorem in a convenient form for lattice systems of interacting particles described by Hamiltonians that are not necessarily translation-invariant with possibly infinite-range pair-potentials and with general complete metric spaces as single-spin spaces. Our reformulation includes existence, and covers both classical lattice sytems and the Euclidean version of quantum lattice systems. Further, a generalization to the Dobrushin-Shlosman theorem for translation-invariant lattice systems is given. Subsequently, we give a series of applications of these uniqueness criteria to some high-temperature classical lattice systems including the Heisenberg, Potts and Ising models. An application to classical lattice systems for which the local Gibbs measures are convex perturbations of Gaussian measures is also given.
@@
The existence problem. This may be solved by constructing Gibbs measures as thermodynamic limits. This may be achieved by proving that the family of local Gibbs measures, indexed by an increasing sequence of bounded regions filling the whole lattice, has at least one limit point in the weak topology, and that this limit point is a Gibbs measure. Sufficient conditions were derived by Dobrushin [19, 20, 22] and [23, Thm. 1]. See also [74, Thm. 1.3]. Dobrushin’s existence criteria have been applied to some classical lattice models of Euclidean lattice field theories with single-spin space $\mathbb{R}$ in [17, 11]. Therein, some restrictions on the boundary conditions are necessary. One of the key methods is Ruelle’s technique of superstability estimates in [68, 69, 50] which requires the interactions to be translation invariant, superstable and with pair-potentials growing at most quadratically. Applied to some models of Euclidean lattice field theories with single-spin space $\mathbb{R}$, it is proved in [17] that the family of local Gibbs measures associated with a wide class of boundary conditions has at least one accumulation point in the set of ’superstable’ Gibbs measures, a subset of the set of tempered Gibbs measures, see [17, Thm. 1.2]. This result was extended to some quantum anharmonic lattice systems with superstable interactions in [60, Thm 2.6] by extending Ruelle’s technique to quantum statistical mechanics, see [57, 58]. See also [46, Thm. 3.1] covering a wide class of quantum anharmonic lattice systems. In general, for quantum lattice systems where the single-spin spaces are infinite-dimensional, verifying Dobrushin’s existence criteria turns out to be challenging, see, e.g., [74, Sec. I.5]. A brief review of methods can be found in [7, Sec. 2].
@@
We mention that Dobrushin’s uniqueness criterion has been applied to some classical lattice models of Euclidean lattice field theories with single-spin space $\mathbb{R}$ in [17, Sec. 2]. The key ingredient is an expression of the Wasserstein distance for probability measures on $\mathbb{R}$ in terms of their distribution functions, see [76, Theorem]. Dobrushin’s uniqueness criterion has also been applied to some quantum anharmonic lattice models of quantum crystals in [3, 5]. See also [46, Thm 3.4]. Therein, the main ingredient is the representation of the Wasserstein distance by the Kantorovich-Rubinstein duality theorem, see, e.g., [30]. For completeness, we mention some other techniques used to prove uniqueness of Gibbs measures. Cluster expansion techniques have been used in [61] to prove uniqueness in the high-temperature regime for a class of quantum anharmonic lattice systems with superstable interactions. Note that in 1-dimensional lattices, uniqueness holds at all temperatures. See also, e.g., [2, 64, 77]. In [4], a quantum crystal with double-well potential is shown to have a phase transition using a Peierls-type argument.
@@
In high-temperature translation-invariant lattice systems with interactions in the Dobrushin uniqueness region, Dobrushin’s uniqueness theorem assures the non-existence of first-order phase transitions. Gross showed in fact the absence of second-order phase transitions. More precisely, Gross proved in [36] that the pressure is twice continuously Gateaux differentiable in the Dobrushin uniqueness region (differentiability is in parameters occurring linearly in the interactions). Prakash extended the Gateaux differentiability to any order in a subset of the Dobrushin uniqueness region, see [63]. Turning to complex analyticity, Gallavotti and Miracle-Sole showed in the case of single-site spin space consisting of two points, that the pressure is analytic at high temperature for a broader class of potentials than that to which the Dobrushin uniqueness theorem applies, see [32]. The techniques used seem to be limited to spin spaces having a finite number of points. For some results on high-temperature analyticity with continuous spin spaces and a general class of many-body long-range interactions, see, e.g., [43]. Herein, Dobrushin’ uniqueness techniques are used. See also [26, 27]. For some general results on analyticity via high-temperature expansions, see, e.g., [29, 42] and references therein.
@@
In this paper, we revisit Dobrushin’s uniqueness theorem of Gibbs measures for lattice systems at thermal equilibrium that are not necessarily translation-invariant, consisting of interacting particles (with possibly long-range interactions) and with a general complete metric space as single-spin space. Both the existence and uniqueness problems are proved in this general setting, along with the Dobrushin-Shlosman generalization of Dobrushin’s uniqueness theorem for translation-invariant lattice systems. Our formulation of the uniqueness theorems covers both classical lattice systems and the Euclidean version of quantum lattice systems.
@@
Our first series of applications includes the classical Heisenberg model, $q$-state Potts model and Ising model. With a view to applying the Dobrushin and Dobrushin-Shlosman criteria, our focus is on different techniques to estimate the Wasserstein distance between two probability measures. We start by proving a general result of uniqueness at high temperatures with direct application to the classical Heisenberg model, see Proposition 6.1. Our estimation of the Wasserstein distance is based on an approximation procedure for probability measures. Subsequently, we derive an explicit lower bound for the critical inverse temperature for the $q$-state Potts models in dimension $d=2,3$ (asymptotic for large $q$), followed by the $d$-dimensional Ising model. The results are based on an explicit formula of the Wasserstein distance for probability measures on $q$-point spaces, see Lemma 6.3. Finally, the effectiveness of the Dobrushin-Shlosman criteria is illustrated with the square lattice Ising model in dimension $d=2,3$. The estimation of the Wasserstein distance is based on linear programming techniques, see, e.g., [33] and Section 6.3.1 for a brief overview. This algorithm is used to numerically compute a lower bound for the critical inverse temperature.
@@
Our second series of applications focuses on convex perturbations of the Gaussian free field model. After applying the Dobrushin criterion to the Gaussian free field model with $d$-dimensional spins, the uniqueness result is extended when including a perturbation by a convex self-interaction. Proposition 7.5 covers the case of 1-dimensonial spins while Proposition 7.9 covers the case of $d$-dimensional spins for a particular class of convex potentials. We show that uniqueness holds at all temperatures by constructing suitable couplings with the 1-point Gibbs measures as marginals.
@@
Our paper is organized as follows. In Section 2, we recall some properties of the Wasserstein distance and give the completeness result in Theorem 2.1. After recalling some definitions of local and limit Gibbs distributions, Theorem 3.5 in Section 3 is our reformulation of Dobrushin’s uniqueness theorem for general lattice systems of interacting particles when the single-spin space is a general metric space. This covers both classical lattice systems and the Euclidean version of quantum lattice systems. Two corollaries related to the decay of correlations in the case of nearest-neighbour interactions are also given. The proof of Theorem 3.5 and proofs of the corollaries are given in Section 4. Theorem 5.1 in Section 5 is our reformulation of Dobrushin-Shlosman’s uniqueness theorem for translation-invariant lattice systems. In Section 6, we focus on a first series of applications of the uniqueness criteria to classical lattice systems at high-temperatures. Dobrushin’s criterion is applied to the classical Heisenberg model in Section 6.1, the $q$-state Potts model in Section 6.2.1 and the $d$-dimensional Ising model in Section 6.2.3. Before applying the Dobrushin-Shlosman criterion to the 2 and 3-dimensional square lattice Ising model in Section 6.3.2, we recall some results from linear programming techniques in Section 6.3.1 including the algorithm used for the numerical estimations. In Section 7, we focus on a second series of applications to classical lattice systems with convex perturbations. Dobrushin’s criterion is applied to the Gaussian free field model with $d$-dimensional spins in Section 7.1. The perturbation by a convex self-interaction potential is then treated in Section 7.2, see Propositions 7.5 and 7.9. Section 8 contains the proof of Theorem 2.1. The key result for the proof of Theorem 2.1 $(ii)$ is Proposition 8.5. Section 10 is our Appendix. The disintegration theorem for measures is recalled in Section 10.1. The dual problem corresponding to the linear programming problem is discussed in Section 10.2. By way of application, a proof of the Kantorovich-Rubinstein duality theorem is given in Section 10.2.3. The quantum harmonic crystal lattice model is revisited in Section 10.3, see Proposition 10.9.
@@
We now turn to the definition of limit Gibbs distributions. We refer the reader to the disintegration theorem in Sec. 10.1 for the existence of conditional probability measures.
@@
The conditional probability measure in $(ii)$ above is relative to the projection map $\Pi_{\Lambda^{c}}$, see Sec. 10.1. Denoting it by $\mu^{\beta}(\cdot\,|\underline{\eta}_{\Lambda^{c}})$ for a fixed configuration $\underline{\eta}_{\Lambda^{c}}\in\mathcal{X}^{\Lambda^{c}}$, $(ii)$ reads
@@
In Theorem 3.5, $(\mathcal{X},\rho)$ is a general complete metric space. The latter could be the infinite-dimensional space of all continuous and periodic functions on $[0,\beta]$ equipped with the supremum norm, so that Theorem 3.5 also applies to quantum lattice systems. See Sec. 10.3.
@@
Clearly, the first term in the r.h.s. of (4.17) tends to 0 in the limit $p\rightarrow\infty$. To conclude the proof of (4.14), we use the monotone convergence theorem to take $n\rightarrow\infty$.
@@ 5 A generalization: The Dobrushin-Shlosman Theorem
$r$ in (5.1)-(5.3) may depend on the boundary condition $\eta_{\Lambda_{0}^{c}}$.
@@
Assume now that $\epsilon<\frac{\alpha}{2}$, see (5.1), and choose $\Lambda^{\prime}\subset\mathbb{Z}^{d}$ such that $\Lambda+\Lambda_{0}\subset\Lambda^{\prime}$ and so large that
$R\sum_{k\in\Lambda^{\prime c}}\mathrm{e}^{-\frac{\alpha}{2}|k|}<\frac{1}{2}(1-\lambda).$ (5.12)
@@
Here, we used (5.1) and (5.12) in the right-hand side of the first and second inequality respectively. From (5.10) together with (5.11) and (5.13), we eventually have
@@
The proof crucially relies on the condition (5.1), corresponding to short-range interactions, in obtaining the bound (5.13). If the interactions are finite-range, the condition (5.1) is superfluous.
Remark 5.5
The proof easily extends to tempered Gibbs measures. Alternatively, the exponential decay in (5.1) can be replaced by a sufficiently fast polynomial decay provided that the distributions satisfy the condition in (3.11).
@@
[…]
@@
Further, let $\mu_{0}\in\mathcal{P}(\mathcal{X})$ be a given a priori probability measure and, for a given $\beta>0$, let $\mu_{j}^{\beta}=\mu^{\beta}$ denote the associated 1-point Gibbs distribution generated by (6.1). Let $C_{1}>0$ be defined as
@@
In fact, considering $\mathbb{S}^{r}$ as a subset of $\mathbb{R}^{r+1}$, the estimate can be slightly improved to $\beta<(24dJ)^{-1}$. Using the symmetry as in the Ising model in Sec. 6.2.3 below, the estimate can be improved further.
@@
followed by assumption $(ii)$, we obtain, under the conditions (6.7), that
@@
$\mu^{\beta}(A|\underline{\eta}^{\prime})-\mu^{\beta}(A|\underline{\eta})=\\ \frac{1}{Z^{\beta}(\underline{\eta})}\int_{A}\left(\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta}^{\prime})-\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta})\right)\mu_{0}(d\xi_{0})-\frac{\mu^{\beta}(A|\underline{\eta}^{\prime})}{Z^{\beta}(\underline{\eta})}\int_{\mathcal{X}}\left(\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta}^{\prime})-\mathpzc{f}^{\beta}(\xi_{0}|\underline{\eta})\right)\mu_{0}(d\xi_{0}).$
@@
Note that $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X}\times F)=\mu^{\beta}(F|\underline{\eta}^{\prime})$ and $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(E\times\mathcal{X})=\mu^{\beta}(E|\underline{\eta})$ so that $\sigma_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}\in\Xi_{\mathcal{X}}(\mu^{\beta}(\cdot\,|\underline{\eta}),\mu^{\beta}(\cdot\,|\underline{\eta}^{\prime}))$. This follows from the fact that $\nu_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X})=\tilde{\nu}_{\underline{\eta},\underline{\eta}^{\prime}}^{\beta}(\mathcal{X})=\frac{2\kappa_{\beta}}{1+\kappa_{\beta}}$. Now,
@@
The $q$-state Potts model is a generalization of the Ising model in that the spin values can be any arbitrary integer in $\{1,\ldots,q\}$, see, e.g., [62]. We refer to, e.g., [10, 39, 78], for some exact results in dimension $d=2$.
@@
Below, this expression is evaluated numerically in dimensions $d=2$ and $d=3$. When $d=2$, there are in general only 9 distinct cases to be considered. Namely, since the interaction is invariant under renaming of spins, the possible distinct configurations for $\underline{s}$ are
$(1,1,1,1),\quad(1,1,1,2),\quad(1,1,2,2),\quad(1,1,2,3),\quad(1,2,3,4).$
Changing one spin in the first configuration, it can only combine with the second, i.e. $\underline{s}^{\prime}=(1,1,1,2)$. If $\underline{s}=(1,1,1,2)$ then, changing $2$ back to $1$, results in nothing new. Changing $2$ to $3$, however, we obtain $\underline{s}=(1,1,1,3)$, which, while equivalent thermodynamically, has a non-zero Wasserstein distance to $(1,1,1,2)$. Changing one of the spins $1$ to $2$ leads to $\underline{s}^{\prime}=(1,1,2,2)$, and changing it to a different value leads to $\underline{s}^{\prime}=(1,1,2,3)$. If $\underline{s}=(1,1,2,2)$ then the spins equal 1 and 2 are equivalent. We therefore only need to consider changing a 2-spin. Changing it back to a 1, leads to a previous case (with $\underline{s}$ and $\underline{s}^{\prime}$ exchanged), whereas changing it to a 3 gives $\underline{s}^{\prime}=(1,1,2,3)$. If $\underline{s}=(1,1,2,3)$ then the spins equal 2 and 3 are equivalent. Changing the 3 to a 1 or 2 leads to a previous combination. Changing 3 to a 4 leads to $\underline{s}^{\prime}=(1,1,2,4)$, which as in the case $(1,1,1,2)$ with $(1,1,1,3)$ has a non-zero Wasserstein distance to $\underline{s}$. Alternatively, we can change a 1 to a 2, leading to $\underline{s}^{\prime}=(1,2,2,3)$ or we can change a 1 to another spin value, leading to $\underline{s}^{\prime}=(1,2,3,4)$. Finally, if $\underline{s}=(1,2,3,4)$, we can replace one of these spins by another value, say 5, leading to $\underline{s}^{\prime}=(1,2,3,5)$. In all, we have the following possible pairings:
@@
$\underline{s}$ $=$ $(1,1,2,3),\quad\underline{s}^{\prime}=(1,2,2,3)$
$\underline{s}$ $=$ $(1,1,2,3),\quad\underline{s}^{\prime}=(1,2,3,4)$
$\underline{s}$ $=$ $(1,2,3,4),\quad\underline{s}^{\prime}=(1,2,3,5).$
Of course, if $q<5$ then some of these combinations are excluded. Figures 1 and 2 below show the graphs of $\rho_{W}$ as a function of $\beta$ for these cases with $q=4$ and $q=30$ respectively.
@@
In Figures 1 and 2, the line $y=\frac{1}{4}$ has also been drawn, which is the upper bound on the right-hand side of the Dobrushin inequality (3.15). For $q=4$, the curve which intersects the horizontal line in the left-most point corresponds to the case $\underline{s}=(1,1,1,2)$ and $\underline{s}^{\prime}=(1,1,2,2)$. For $q=30$, this curve is overtaken by the case $\underline{s}=(1,1,1,1)$ and $\underline{s}^{\prime}=(1,1,1,2)$. Note that the corresponding curve for $q=4$ is the blue curve with a clear maximum. The change-over occurs at $q=20$, as is clearly visible in Figure 3 below showing the Dobrushin lower bound for $\beta_{c}J$ ($\beta_{c}$ is the critical inverse temperature) as a function of $q$.